OpenStudy (anonymous):

help fast please! The midpoint of a line segment with end points as (x, 4) and (-2, -2) is (2, 1). What is the value of x? x=6 x=4 x=-3 x=-2 I know how to find the midpoint when given two points but this already gives the midpoint...:(

jimthompson5910 (jim_thompson5910):

to find the midpoint, you average the coordinates ie you add up the corresponding coordinates and divide by 2 so.... (x1+x2)/2 = x coordinate of midpoint (x+2)/2 = -2 ... plug in the given values now solve (x+2)/2 = -2 for x to get your answer

OpenStudy (anonymous):

you are using the Midpoint formula and basically substituting the value for the Midpoint calculated as well to figure out the x.

OpenStudy (anonymous):

including the two points given.

OpenStudy (anonymous):

so (-3 + 2) / 2 = -2 right?

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

(x+2)/2 = -2 x+2 = -2*2 x+2 = -4 x = ???

OpenStudy (anonymous):

i think... x + 2 = - 4 - 2 -2 x = -6 ???

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

i'm confused -6 isn't one of the answer choices...

jimthompson5910 (jim_thompson5910):

then there must be a typo somewhere

jimthompson5910 (jim_thompson5910):

because you have the correct answer

OpenStudy (anonymous):

Let me check, give me a second...

OpenStudy (anonymous):

The midpoint of a line segment with end points as (x, 4) and (-2, -2) is (2, 1). What is the value of x? x = 6 x = 4 x = -3 x = -2 This is a copy and paste of the question, there is no typo.... :(

jimthompson5910 (jim_thompson5910):

if x = -6, then... (-6+2)/2 -4/2 -2 So adding up the x coordinates -6 and 2 and dividing by 2 gives -2, which is the x coordinate of the midpoint

jimthompson5910 (jim_thompson5910):

so either there's a typo in the choices or the original points

OpenStudy (anonymous):

idk can't find one...

jimthompson5910 (jim_thompson5910):

say that the endpoint wasn't (2,1) and instead it was (-2, 1) (x-2)/2 = -2 x-2 = -2*2 x - 2 = -4 x = -4 + 2 x = -2 So the answer would be x = -2 if the endpoint was (-2, 1) but again this is just a guess

jimthompson5910 (jim_thompson5910):

well not a typo that's your fault, i meant a typo by your teacher or the book's author

OpenStudy (anonymous):

So there no other than to just pick one answer randomly right? i'll just pick A because its closest to what our conclusion was...

jimthompson5910 (jim_thompson5910):

oh wait, i completely mixed up the endpoint and midpoint because i read it out of order my bad the midpoint is (2,1) the endpoints are (x, 4) and (-2, -2) (x-2)/2 = 2 x-2 = 2*2 x-2 = 4 x = 4+2 x = 6

jimthompson5910 (jim_thompson5910):

so the answer is actually 6

jimthompson5910 (jim_thompson5910):

gotta read the instructions more carefully lol

OpenStudy (anonymous):

hehe :) well thanks for everything and now i know what equation to use if i get a question like this again ;)

jimthompson5910 (jim_thompson5910):

you're welcome